Can you please help me solve this problem. Thank you!

Determine whether the New York Yankees or New York Mets had a relatively better home run hitting season. The yankees hit 205 home runs and play in the American League, where the mean number of home runs was 192. The standard deviation was 36.0. The Mets hit 198 home runs and play in the National league, where the mean number of home runs was 186.9. The standard deviation was 28.7

Answer by stanbon(75887)

Determine whether the New York Yankees or New York Mets had a relatively better home run hitting season.
The yankees hit 205 home runs and play in the American League, where the mean number of home runs was 192. The standard deviation was 36.0.
--------
z(205) = (205-192)/36 = 13/36 = 0.3611
----------------------------
The Mets hit 198 home runs and play in the National league, where the mean number of home runs was 186.9. The standard deviation was 28.7
z(198) = (198-186.9)/28.7 = 0.3862
===
Since the Met's z-score is slightly higher than the Yankees
the Mets had a slightly better home run hitting season.
===========
Cheers,
Stan H.